OpenStudy (anonymous):

What is the solution of square root 5x+1 - square root x = 5?

OpenStudy (phi):

\[ \sqrt{5x+1} - \sqrt{x} = 5 \]?

OpenStudy (anonymous):

yes

OpenStudy (phi):

add - sqrt(x) to both sides to get \[ \sqrt{5x+1} = 5 +\sqrt{x} \] now square both sides. can you do that ?

OpenStudy (phi):

*add +sqrt(x) to both sides

OpenStudy (anonymous):

5x+1=25+x

OpenStudy (phi):

the 5x+1 is on the left side is ok. but not the right side. do you know FOIL? (5 + sqr(x)) ( 5 + sqr(x) ) gives 4 terms. try again

OpenStudy (phi):

you will get \[ 25 + 5 \sqrt{x} + 5 \sqrt{x} +x \] you can combine the "like terms" to simplify it to \[ 25 + 10 \sqrt{x} +x \] and your problem is \[ 5x+1=25 + 10 \sqrt{x} +x \] now add -25 to both sides and -x to both sides. can you do that ?